The ratio of radii of first shell of H atom and that of fourth shell of He ion is 
1) 1:8
2) 1:4
3) 1:√8
4) 1:3 
 
Correct ans is option 1 plis explain how

Grade:12th pass

1 Answers

Khimraj
5 years ago
Radius = 0.529 n²/Z Å
R1/R2 = n12*Z2/n22*Z1 = 12*2/42*1 = 1:8
Hope it clears.............................................